8. Problems and Applications Q2 American and Japanese workers can each produce 4 cars a year. An American worker can produce 10 tons of grain a year, whereas a Japanese worker can produce 5 tons of grain a year. To keep things simple, assume that e ach country has 100 million workers. Complete the following table with the number of workers needed to make one car or 1 ton of grain in the United States and Japan. Workers Needed to Make 1 Car 1 Ton of Grain United States 4 10Y Use the blue line (circle symbol) to graph the production possibilities frontier for the American economy. Then use the green line (triangle symbol) to graph the production possibilities frontier for the Japanese economy U.S Japan 150
graph the production possibilities frontier for the Japanese economy U.S o 100 200 300 400 00600 T00001000 ty cost of a car and of a ton of grain for both the United States and Japan e of (In terms of tons of up) (In terms of cars given up) 2 1/2 Y 2/5 Japan 11/4Y Y has an absolute advantage Given this information, in producing grain.
Opportunity Cost of 1114 ▼ Given this information, in producing grain. (Millions) OTrue
Answer #1
cars 1 ton of grain
United States .25 .10
Japan .25 .20

United States =1 worker can produce 4 cars

so 1 car can be produced by .25 workers

similarly, 1 worker can produce 10 tons of grain

so 1ton of grain can be produced by .10 workers

In Japan,

1 worker can produce 4 cars

so 1 car can be produced by .25 workers

similarly, 1 worker can produce 5 tons of grain

so 1ton of grain can be produced by .20 workers

production possibility curve has been correctly drawn in the questions

I.e in United States , 1 worker = 4 cars

100 million workers = 400 million cars

or

1 worker = 10 tons of grain

100 million workers = 1000 million tons of grain

in Japan , 1 worker = 4 cars

100 million workers = 400 million cars

or

1 worker = 5 tons of grain

100 million workers = 500 million tons of grain

opportunity cost table

1 car 1 ton grain
United States 2.5= .25/.10 0.4= .10/.25
Japan 1.25 = .25/.20 0.8= .20/.25

Given this information , Japan/United States has an absolute advantage in producing cars while United States has an absolute advantage to produce 1 tons of grain ( as workers required to produce 1 ton grain is less in United States .)

Also , Japan has an comparative advantage in producing cars(as opportunity cost is lower in Japan) while United States has an comparative advantage to produce 1 tons of grain( as opportunity cost is lower in United States)

if workers divided half in each good production then

cars (million ) grain (million )
United States 200 = 50*4 500 = 50 *10
Japan 200= 50*4 250 = 50*5

true , both nations would be better off if they produced the good in which they specialise and then trade 200million tons of grain with 100 million cars .